You charge a piece of aluminum foil (mass = 7.55 g) by touching it to a charged rod. The charged rod gives the aluminum foil a charge of 10
Answer #1

Here ,

let the magnitude of electric field is E ,

Now, for aluminium plate to hover,

mg = q*E

0.00755 * 9.8 = 10 *10^-6 * E

E = 7399 N/C

the value of the electric field from the charged plate is 7399 N/C